OpenStudy (anonymous):

write each quotient as a rational expression and simplify 1. 7 divided by 3/5 2. - 8/7 divided by - 4/3 3. 3ab/4cd divided by 9a/8d

OpenStudy (anonymous):

\[\frac{7}{\frac{3}{5}}=7\times \frac{5}{3}=\frac{35}{3}\] \[\frac{-\frac{8}{7}}{-\frac{4}{3}}=\frac{8}{7}\times \frac{3}{4}=\frac{6}{7}\]

OpenStudy (anonymous):

and finally \[\frac{3ab}{4bc}\div \frac{9a}{8d}=\frac{3ab}{4bc}\times\frac{8d}{9a}=\frac{2d}{c}\]

OpenStudy (anonymous):

Thank you.